### v1

# Degenerate scalar scenario of two Higgs doublet model with a complex singlet scalar

Abstract
We study the two Higgs doublet model with a complex singlet scalar whose imaginary part acts as a dark matter (DM). The scattering of DM and quarks, mediated by three CP-even scalars in this model, is suppressed when masses of CP-even scalars are degenerate; that is called the “degenerate scalar scenario”. Based on this scenario, we show that the strong first-order electroweak phase transition (EWPT) can be achieved without conflicting with constraints from the DM relic density and the direct detection experiments. We also discuss a shift of scalar trilinear coupling from the Standard Model prediction, which could be a test of this model in collider experiments.
preprint: OCHA-PP-380
I Introduction
Dark matter (DM), whose existence has been suggested by cosmological observations, predicts physics beyond the Standard Model (SM). Among the various DM candidates, weakly interacting massive particles (WIMP) are attractive because their abundance can be explained thermally. (the so-called “WIMP miracle”). However, despite the vigorous search, experiments such as accelerator experiments and DM direct detection experiments have not found the DM signal. In particular, the LZ experiment Aalbers et al. (2022) places very strong constraints on the scattering cross sections of WIMP-DM and nucleons, posing a major challenge for models that include WIMP-DM [^1] The updated results of the LZ experiment, which were presented at the international conference “TeV Particle Astrophysics” held at the University of Chicago in August 2024, impose even stricter constraints..
Possible scenarios consistent with the current direct detection experiments include (1) the mass and interaction of WIMP-DM are decoupled from the SM, and (2) DM-nucleon scattering is suppressed by a built-in mechanism in the model. Since no new particles have been found after the Higgs boson discovery, focusing on the second scenario is natural. One model that realizes this scenario is the complex singlet scalar extension of the Standard Model (CxSM) Barger et al. (2009), which includes a pseudo Nambu-Goldstone (pNG) DM. In the CxSM, as the name suggests, the SM is extended by adding a complex scalar single particle field $S$. The imaginary part of $S$ behaves as the pNG DM, whose stability is guaranteed by the CP symmetry of the scalar potential. On the other hand, the real part of $S$ mixes with the SM Higgs boson to form the mass eigenstates $h_{1}$ and $h_{2}$. These two particles mediate DM-quark scattering. In ref. Abe et al. (2021), it is proposed that when the masses of the two Higgs bosons are degenerate, i.e., $m_{h_{1}}\simeq m_{h_{2}}$, this scattering is suppressed. This suppression mechanism is called the degenerate scalar scenario.
On the other hand, baryon asymmetry of the Universe (BAU), along with DM, is one of the unsolved problems. The most testable scenario to realize BAU is electroweak baryogenesis (EWBG) Kuzmin et al. (1985); Rubakov and Shaposhnikov (1996); Funakubo (1996); Riotto (1998); Trodden (1999); Bernreuther (2002); Cline (2006); Morrissey and Ramsey-Musolf (2012); Konstandin (2013); Senaha (2020) associated with the Higgs physics. EWBG requires strong first-order electroweak phase transition (EWPT), which suggests the need for an extension of the SM Kajantie et al. (1996); Rummukainen et al. (1998); Csikor et al. (1999); Aoki et al. (1999). The CxSM is an attractive model for both DM and BAU, but in ref. Cho et al. (2021), the poor compatibility between the degenerate scalar scenario and strong first-order EWPT driven by the tree-level structure of the effective potential has been pointed out. In order to explain them simultaneously, the DM mass needs to be half the Higgs mass, but here the DM relic abundance is very small due to the resonance effect of the DM annihilation process, and is far from the observed value $\Omega_{\mathrm{DM}}h^{2}=0.1200\pm 0.0012$ Aghanim et al. (2020).
In this paper, we consider a model in which a SU(2)L scalar doublet and a SU(2)L complex scalar singlet are added to the SM. Since this model is the two Higgs doublet model (2HDM) with the complex singlet $S$, we refer to this model as 2HDMS. In the 2HDMS, the CP-even scalars of the two Higgs doublet and the real part of $S$ mix to form three Higgs particles. The imaginary part of $S$ acts as the DM. We investigate the feasibility of the degenerate scalar scenario for four types of Yukawa interaction between the SM fermions and the SU(2)L doublet Higgs bosons, which are distinguished by the $Z_{2}$ charge of each fermion. In the 2HDMS, the one-loop effect is most important for strong first-order EWPT, unlike the CxSM, where the tree-level structure is extremely important. As a result, it is shown that the degenerate scalar scenario is sufficiently valid even in the parameter region where strong first-order EWPT is achieved. In order to resolve the conflict between the degenerate scalar scenario and strong first-order EWPT, a two-component DM model that adds a fermion DM to the CxSM is also being studied Cho et al. (2024), but here we explain the current DM abundance using only the pNG DM of the 2HDMS.
The structure of this paper is as follows. Sec. II gives a brief review of the 2HDMS. After introducing the degenerate scalar scenario in Sec. III, to discuss the compatibility of the degenerate scalar scenario and strong first-order EWPT, we qualitatively analyze EWPT in Sec. IV. Then, we numerically investigate the parameter region consistent from the perspective of DM and EWPT in Sec. V. Lastly, Sec. VI summarizes this study.
II The Model
In the 2HDMS, the scalar potential is given as
$V_{0}=V_{0,\mathrm{2HDM}}+V_{0,\mathrm{S}}.$ (1)
The first term in r.h.s. in (1) contains only the doublet Higgs fields $\Phi_{1}$ and $\Phi_{2}$
[…]
$+\frac{\lambda_{1}}{2}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)^{2}+\frac{\lambda_{2}}{2}\left(\Phi_{2}^{\dagger}\Phi_{2}\right)^{2}+\lambda_{3}\left(\Phi_{1}^{\dagger}\Phi_{1}\right)\left(\Phi_{2}^{\dagger}\Phi_{2}\right)$
[…]
where only $m_{3}^{2}$ term breaks the $Z_{2}$ symmetry of the doublets softly. The second term in (1) is given by the singlet $S$ and doublets as
$V_{0,\mathrm{S}}\left(\Phi_{1},\Phi_{2},S\right)$ $=\frac{\delta_{1}}{2}\Phi_{1}^{\dagger}\Phi_{1}|S|^{2}+\frac{\delta_{2}}{2}\Phi_{2}^{\dagger}\Phi_{2}|S|^{2}+\frac{b_{2}}{2}|S|^{2}+\frac{d_{2}}{4}|S|^{4}$
[…]
where the terms in the first line are invariant under a global U(1) transformation $\quantity(S\to e^{i\theta}S)$, and terms in the second line break the symmetry softly. In the following, the scalar potential is assumed to be CP-invariant, i.e., all coefficients are real. Three scalar fields $\Phi_{i}~{}(i=1,2)$ and $S$ can be written as
$\Phi_{i}=\binom{\phi_{i}^{+}}{\frac{1}{\sqrt{2}}\left(v_{i}+h_{i}+i\eta_{i}\right)},\quad S=\frac{1}{\sqrt{2}}\left(v_{S}+s+i\chi\right).$ (4)
At this time, $v=\sqrt{v_{1}^{2}+v_{2}^{2}}$ = 246.22 GeV and $\tan\beta=v_{2}/v_{1}$ is defined. The following unphysical scalar particles are present in this model; the charged scalar $\quantity(\phi_{i}^{+})$, the CP-odd scalar $\quantity(\eta_{i})$, and the CP-even scalars $\quantity(h_{1},h_{2},s)$. On the other hand, a pseudoscalar $\chi$ is the physical DM candidate.
First derivatives of $V_{0}$ with respect to $h_{1},h_{2}$ and $s$ are respectively given by
$\left\langle\frac{\partial V_{0}}{\partial h_{1}}\right\rangle$ $=m_{1}^{2}v_{1}-m_{3}^{2}v_{2}+\frac{\lambda_{1}}{2}v_{1}^{3}+\frac{\lambda_{345}}{2}v_{1}v_{2}^{2}+\frac{\delta_{1}}{4}v_{1}v_{S}^{2}=0,$ (5)
$\left\langle\frac{\partial V_{0}}{\partial h_{2}}\right\rangle$ $=m_{2}^{2}v_{2}-m_{3}^{2}v_{1}+\frac{\lambda_{2}}{2}v_{2}^{3}+\frac{\lambda_{345}}{2}v_{1}^{2}v_{2}+\frac{\delta_{2}}{4}v_{2}v_{S}^{2}=0,$ (6)
$\left\langle\frac{\partial V_{0}}{\partial s}\right\rangle$ $=\sqrt{2}a_{1}+\frac{b_{1}+b_{2}}{2}v_{S}+\frac{\delta_{1}}{4}v_{1}^{2}v_{S}+\frac{\delta_{2}}{4}v_{2}^{2}v_{S}+\frac{d_{2}}{4}v_{S}^{3}=0,$ (7)
where $\lambda_{345}=\lambda_{3}+\lambda_{4}+\lambda_{5}$. The tree-level masses of the CP-even scalars are obtained by
$-\mathcal{L}_{\text{mass}}$ $=\frac{1}{2}\left(\begin{array}[]{lll}h_{1}&h_{2}&s\end{array}\right)\mathcal{M}_{S}^{2}\left(\begin{array}[]{c}h_{1}\\ h_{2}\\ s\end{array}\right)=\frac{1}{2}\left(\begin{array}[]{lll}H_{1}&H_{2}&H_{3}\end{array}\right)O^{\top}\mathcal{M}_{S}^{2}O\left(\begin{array}[]{c}H_{1}\\ H_{2}\\ H_{3}\end{array}\right)$ (16)
$=\frac{1}{2}\sum_{i=1}^{3}m_{H_{i}}^{2}H_{i}^{2},$ (17)
with
$\mathcal{M}_{S}^{2}=\left(\begin{array}[]{ccc}m_{3}^{2}\frac{v_{2}}{v_{1}}+\lambda_{1}v_{1}^{2}&-m_{3}^{2}+\lambda_{345}v_{1}v_{2}&\frac{\delta_{1}}{2}v_{1}v_{S}\\ -m_{3}^{2}+\lambda_{345}v_{1}v_{2}&m_{3}^{2}\frac{v_{1}}{v_{2}}+\lambda_{2}v_{2}^{2}&\frac{\delta_{2}}{2}v_{2}v_{S}\\ \frac{\delta_{1}}{2}v_{1}v_{S}&\frac{\delta_{2}}{2}v_{2}v_{S}&-\frac{\sqrt{2}a_{1}}{v_{S}}+\frac{d_{2}}{2}v_{S}^{2}\end{array}\right),$ (21)
The mixing matrix $O$ is parametrized as
$O(\alpha_{i})$ $=\left(\begin{array}[]{ccc}1&0&0\\ 0&c_{3}&-s_{3}\\ 0&s_{3}&c_{3}\end{array}\right)\left(\begin{array}[]{ccc}c_{2}&0&-s_{2}\\ 0&1&0\\ s_{2}&0&c_{2}\end{array}\right)\left(\begin{array}[]{ccc}c_{1}&-s_{1}&0\\ s_{1}&c_{1}&0\\ 0&0&1\end{array}\right),$ (31)
where $s_{i}=\sin\alpha_{i}$ and $c_{i}=\cos\alpha_{i}~{}(i=1,2,3)$. Note that the mixing matrix is orthogonal, i.e.,
$\sum_{k}O_{ik}O_{jk}=\delta_{ij}.$ (32)
Three Higgs particles appear in the 2HDMS, and $H_{1}$ is the Higgs boson observed at the LHC experiment.
On the other hand, the mass eigenstates of the charged scalar $\phi^{+}$ and the CP-odd scalar $\eta_{i}$ are defined as follows, respectively:
$\quad\binom{\phi_{1}^{+}}{\phi_{2}^{+}}=R(\beta)\binom{G^{+}}{H^{+}},\quad\binom{\eta_{1}}{\eta_{2}}=R(\beta)\binom{G^{0}}{A},$ (33)
where
$R(\beta)=\left(\begin{array}[]{cc}\cos\beta&-\sin\beta\\ \sin\beta&\cos\beta\end{array}\right).$ (36)
and $G^{\pm}$ and $G^{0}$ are Nambu-Goldstone bosons. The mass terms of the charged Higgs $H^{\pm}$ and the CP-odd scalars $m_{A}$ are
$m_{H^{\pm}}^{2}$ $=\frac{m_{3}^{2}}{\sin\beta\cos\beta}-\frac{1}{2}\left(\lambda_{4}+\lambda_{5}\right)v^{2},$ (37)
$m_{A}^{2}$ $=\frac{m_{3}^{2}}{\sin\beta\cos\beta}-\lambda_{5}v^{2}.$ (38)
Also, DM mass is given by
$m_{\chi}^{2}=-\frac{\sqrt{2}a_{1}}{v_{S}}-b_{1}.$ (39)
Let us summarize our input parameters. There are 14 degrees of freedom in the scalar potential: $\{m_{1}^{2},m_{2}^{2},m_{3}^{2},\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4},\lambda_{5},\delta_{1},\delta_{2},b_{2},d_{2},a_{1},b_{1}\}$. First, $\{m_{3}^{2},a_{1}\}$ remain as input parameters. $\{m_{1}^{2},m_{2}^{2},b_{2}\}$ are determined by the tadpole conditions (5)-(7) and $\{\lambda_{4},\lambda_{5},b_{1}\}$ are fixed by the particle masses (37),(38) and (39). The other six Lagrangian parameters are determined by the mass matrix of the CP-even scalars. In Appendix A, we list relationships between the input parameters and original Lagrangian parameters.
We mention the tree-level flavor changing neutral current (FCNC) restricted experimentally. The Lagrangian of the Yukawa interaction is given as follows:
$\mathcal{L}_{\text{Yukawa }}^{2\mathrm{HDMS}}=-\bar{Q}_{L}Y_{u}\tilde{\Phi}_{u}u_{R}-\bar{Q}_{L}Y_{d}\Phi_{d}d_{R}-\bar{L}_{L}Y_{\ell}\Phi_{\ell}\ell_{R}+\mathrm{h.c.},$ (40)
where $\tilde{\Phi}_{u}\equiv i\sigma_{2}\Phi^{*}_{u}$ and $\sigma_{2}$ being the Pauli matrix. ${Q}_{L}$ and ${L}_{L}$ represent left-handed quarks and left-handed leptons, i.e., $Q_{L}=\left(u_{L},d_{L}\right)^{\top},~{}L_{L}=\left(\nu_{L},e_{L}\right)^{\top}$. $u_{R},~{}d_{R}$, and $l_{R}$ are right-handed up-type quarks, right-handed down-type quarks, and right-handed leptons. $Y_{f}(f=u,d,l)$ is the 3$\times$3 Yukawa matrix of the fermions. Here, in order to prevent the tree-level FCNC, we consider the case where one of the doublets couples with each fermion. As shown in Table 1, by assigning the $Z_{2}$ charge for the fermions, we can determine the combination of ${\Phi}_{f}~{}(f=u,d,l)$ and classify them into four types.
$Q_{L},L_{L}$ $u_{R}$ $d_{R}$ $\ell_{R}$ $\Phi_{u}$ $\Phi_{d}$ $\Phi_{l}$
Type-I $+$ -- -- -- $\Phi_{2}$ $\Phi_{2}$ $\Phi_{2}$
Type-II $+$ -- $+$ $+$ $\Phi_{2}$ $\Phi_{1}$ $\Phi_{1}$
Type-X $+$ -- -- $+$ $\Phi_{2}$ $\Phi_{2}$ $\Phi_{1}$
Type-Y $+$ -- $+$ -- $\Phi_{2}$ $\Phi_{1}$ $\Phi_{2}$
Table 1: Assignment of $Z_{2}$ charge to the fermions and combination of the doublets that couples to each fermion.
III Degenerate scalar scenario
We study a suppression mechanism of the scattering process of DM $\chi$ off a quark $q$
$\chi(p_{1})+q(p_{2})\to\chi(p_{3})+q(p_{4})$ (41)
in the 2HDMS. The scattering amplitude $\mathcal{M}$ is given by a sum of three amplitudes $\mathcal{M}_{1},\mathcal{M}_{2}$ and $\mathcal{M}_{3}$ mediated by $H_{1},H_{2}$ and $H_{3}$, respectively:
$i\mathcal{M}$ $=i\left(\mathcal{M}_{1}+\mathcal{M}_{2}+\mathcal{M}_{3}\right),$ (42)
$i\mathcal{M}_{1}$ $=-i2C_{\chi\chi H_{1}}C_{qqH_{1}}\frac{1}{t-m_{H_{1}}^{2}}\bar{u}(p_{4})u(p_{2}),$ (43)
$i\mathcal{M}_{2}$ $=-i2C_{\chi\chi H_{2}}C_{qqH_{2}}\frac{1}{t-m_{H_{2}}^{2}}\bar{u}(p_{4})u(p_{2}),$ (44)
$i\mathcal{M}_{3}$ $=-i2C_{\chi\chi H_{3}}C_{qqH_{3}}\frac{1}{t-m_{H_{3}}^{2}}\bar{u}(p_{4})u(p_{2}),$ (45)
where $t$ describes a momentum transfer, $t\equiv\left(p_{1}-p_{3}\right)^{2}$, and $u(p)~{}(\bar{u}(p))$ represents an incoming (outgoing) quark spinor with a momentum $p$. A factor 2 in the right-hand side of (43), (44) and (45) is a symmetry factor for the $\chi\chi H_{i}$ vertex.
Note that the doublets that couple to each fermion differ depending on the types in the 2HDMS. Therefore, we investigate the suppression mechanism for (A) and (B):
[…]
III.1 Type-I, Type-X
The Yukawa interaction for Type-I and Type-X is given by
$-\mathcal{L}_{\mathrm{Yukawa}}$ $=\sum_{i=1}^{3}C_{uuH_{i}}\overline{u_{L}}u_{R}H_{i}+\sum_{i=1}^{3}C_{ddH_{i}}\overline{d_{L}}d_{R}H_{i},$ (46)
where
$C_{uuH_{i}}=\sum_{i=1}^{3}\frac{m_{u}}{v_{2}}O_{2i},~{}C_{ddH_{i}}=\sum_{i=1}^{3}\frac{m_{d}}{v_{2}}O_{2i},$ (47)
with $m_{u,d}$ being masses of up-type and down-type quarks. On the other hand, the scalar trilinear interaction between $\chi$ and $H_{i}$ is described by
$-\mathcal{L}$ $\supset C_{\chi\chi h_{1}}h_{1}\chi^{2}+C_{\chi\chi h_{2}}h_{2}\chi^{2}+C_{\chi\chi s}s\chi^{2}$
$=C_{\chi\chi H_{i}}H_{i}\chi^{2},$ (48)
where
$C_{\chi\chi H_{i}}$ $\equiv C_{\chi\chi h_{1}}O_{1i}+C_{\chi\chi h_{2}}O_{2i}+C_{\chi\chi s}O_{3i}.$ (49)
Since the momentum transfer $t$ in the direct detection experiments is very small as compared to the mediator masses, i.e., $t\ll m_{H_{i}}^{2}$, the amplitude (42) can be written by
$i\mathcal{M}$ $=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{\left(m_{u}+m_{d}\right)}{v_{2}}\frac{C_{\chi\chi H_{i}}O_{2i}}{m_{H_{i}}^{2}},$
$=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{\left(m_{u}+m_{d}\right)}{v_{2}}\left(C_{\chi\chi h_{1}}\frac{O_{1i}O_{2i}}{m_{H_{i}}^{2}}+C_{\chi\chi{h_{2}}}\frac{O_{2i}^{2}}{m_{H_{i}}^{2}}+C_{\chi\chi s}\frac{O_{3i}O_{2i}}{m_{H_{i}}^{2}}\right).$ (50)
Therefore, due to the orthogonality of the mixing matrix (32), the amplitude vanishes in the degenerate scalar scenario $(m_{H_{1}}=m_{H_{2}}=m_{H_{3}})$.
III.2 Type-II, Type-Y
The Yukawa interaction for Type-II and Type-Y is given by
$-\mathcal{L}_{\mathrm{Yukawa}}$ $=\sum_{i=1}^{3}C_{uuH_{i}}\overline{u_{L}}u_{R}H_{i}+\sum_{i=1}^{3}C_{ddH_{i}}\overline{d_{L}}d_{R}H_{i},$ (51)
where
$C_{uuH_{i}}=\sum_{i=1}^{3}\frac{m_{u}}{v_{2}}O_{2i},~{}C_{ddH_{i}}=\sum_{i=1}^{3}\frac{m_{d}}{v_{1}}O_{1i}.$ (52)
The amplitude (42) becomes
$i\mathcal{M}$ $=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\left(\frac{m_{u}}{v_{2}}\frac{C_{\chi\chi H_{i}}O_{2i}}{m_{H_{i}}^{2}}+\frac{m_{d}}{v_{1}}\frac{C_{\chi\chi H_{i}}O_{1i}}{m_{H_{i}}^{2}}\right)$
$=2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{u}}{v_{2}}\left(C_{\chi\chi h_{1}}\frac{O_{1i}O_{2i}}{m_{H_{i}}^{2}}+C_{\chi\chi{h_{2}}}\frac{O_{2i}^{2}}{m_{H_{i}}^{2}}+C_{\chi\chi s}\frac{O_{3i}O_{2i}}{m_{H_{i}}^{2}}\right)$
$+2i\bar{u}\left(p_{4}\right)u\left(p_{2}\right)\frac{m_{d}}{v_{1}}\left(C_{\chi\chi h_{1}}\frac{O_{1i}^{2}}{m_{H_{i}}^{2}}+C_{\chi\chi{h_{2}}}\frac{O_{2i}O_{1i}}{m_{H_{i}}^{2}}+C_{\chi\chi s}\frac{O_{3i}O_{1i}}{m_{H_{i}}^{2}}\right).$ (53)
As for case A (Type-I, Type-X), this scattering is suppressed when the three Higgs masses are equal.
We comment on the sum rule for the degenerate scalar scenario taking case A (Type-I, Type-X) as an example. The scattering of DM and quarks for case A (Type-I, Type-X) is as shown in Fig. 1. The left side uses the mass eigenstates, while the right side uses the gauge eigenstates.
Figure 1: Feynman diagrams of the DM-quark scattering in the mass and gauge eigenstates of scalar mediators.
In order for the scalar $s$ to couple with the quarks $q$, the scalar $s$ must be converted into $h_{2}$ through the coupling $C_{hs}$ (the right diagram in Fig. 1). On the other hand, the amplitude mediated by $h_{2}$ is proportional to the coupling $C_{\chi\chi h_{2}}$ (the middle diagram in Fig. 1). Therefore, the amplitudes are canceled only when $C_{hs}$ and $C_{\chi\chi h_{2}}$ satisfy a specific relationship.
The scalar potential in the 2HDMS allows us to rewrite the trilinear couplings $C_{\chi\chi h_{1}},C_{\chi\chi h_{2}}$ and $C_{\chi\chi s}$ by the bilinear couplings $C_{h_{1}s},C_{h_{2}s}$ and $C_{ss}$ as
$C_{\chi\chi h_{1}}$ $=\frac{A}{v_{s}}\left(C_{h_{1}s}+\Delta_{1}\right),~{}~{}~{}C_{\chi\chi h_{2}}=\frac{A}{v_{s}}\left(C_{h_{2}s}+\Delta_{2}\right),~{}~{}~{}C_{\chi\chi s}=\frac{A}{v_{s}}\left(C_{ss}+\Delta_{3}\right),$ (54)
where parameters $A,\Delta_{1},\Delta_{2}$ and $\Delta_{3}$ are given by parameters besides $C_{h_{1}s},C_{h_{2}s}$ and $C_{ss}$ in the scalar potential. Extracting the part that is involved in suppressing the scattering, one gets
$\frac{C_{\chi\chi H_{i}}O_{2i}}{m_{H_{i}}^{2}}$ $=\frac{A}{v_{s}}\left\{(C_{h_{1}s}+\Delta_{1})O_{1i}+(C_{h_{2}s}+\Delta_{2})O_{2i}+(C_{ss}+\Delta_{3})O_{3i}\right\}\frac{O_{2i}}{m_{H_{i}}^{2}},$
$=\frac{A}{v_{s}m_{H_{i}}^{2}}\left(O_{2i}O_{3i}m_{H_{i}}^{2}+\Delta_{1}O_{1_{i}}O_{2_{i}}+\Delta_{2}O_{2_{i}}^{2}+\Delta_{3}O_{3_{i}}O_{2_{i}}\right).$ (55)
Therefore, the amplitude in the degenerate Higgs region vanishes when
$\Delta_{2}=0.$ (56)
We show that the condition (56) holds for the scalar potential $V_{0}$ (1). In the 2HDMS, the couplings $C_{h_{2}s}$ and $C_{\chi\chi h_{2}}$ are related to
$V_{0}$ $\supset\frac{\delta_{2}}{2}\Phi_{2}^{\dagger}\Phi_{2}|S|^{2}$
$\supset C_{\chi\chi h_{2}}h_{2}\chi^{2}+C_{h_{2}s}h_{2}s,$ (57)
where
$C_{\chi\chi h_{2}}$ $=\frac{\delta_{2}}{4}v_{2},$ (58)
$C_{h_{2}s}$ $=\frac{\delta_{2}}{2}v_{2}v_{s}.$ (59)
Comparing (58) and (59) with (54), we find
$A$ $=\frac{1}{2},~{}~{}~{}\Delta_{2}=0.$ (60)
Therefore, in the degenerate limit of the mediator masses, the DM-quark scattering is suppressed as long as the potential (1) is used. In other words, the mixing term between $\Phi_{2}$ and $S$ is constrained by the degenerate scalar scenario. If mixing terms such as $\Phi_{2}^{\dagger}\Phi_{2}S,\Phi_{2}^{\dagger}\Phi_{2}S^{2}$ are added, the suppression mechanism does not work [^2] A more detailed analysis, including one-loop scattering, is given in Ref. Cho and Idegawa (2023). This discussion can be applied almost directly to the case of the 2HDMS..
Before finishing this section, we mention the usefulness of the degenerate scalar scenario in the Higgs search experiments at the LHC. In the 2HDMS, the couplings between the Higgs boson $H_{i}$ and the SM particle $X$ are given by the SM couplings multiplied by the mixing matrix elements $O_{1i}$. Taking the decay widths of $H_{i}$ as an example, we obtain
$\Gamma_{H_{1}\rightarrow XX}=\Gamma_{h\rightarrow XX}^{\mathrm{SM}}\left(m_{H_{1}}\right)\times O_{11}^{2},$ (61)
$\Gamma_{H_{2}\rightarrow XX}=\Gamma_{h\rightarrow XX}^{\mathrm{SM}}\left(m_{H_{2}}\right)\times O_{12}^{2},$ (62)
$\Gamma_{H_{3}\rightarrow XX}=\Gamma_{h\rightarrow XX}^{\mathrm{SM}}\left(m_{H_{3}}\right)\times O_{13}^{2},$ (63)
where $\Gamma_{h\rightarrow XX}^{\mathrm{SM}}\left(m_{H_{i}}\right)$ is the decay width of the SM Higgs $h$ as a function of $m_{H_{i}}$. When we add up all the decay processes,
$\Gamma_{H_{1}\rightarrow XX}+\Gamma_{H_{2}\rightarrow XX}+\Gamma_{H_{3}\rightarrow XX}\simeq\Gamma_{h\rightarrow XX}^{\mathrm{SM}}\left(m_{h}\right)$ (64)
is obtained. This SM-like process is effective for all mixing angles, and $H_{i}$ cannot be distinguished in the current experiment Abe et al. (2021).
IV Qualitative discussion for electroweak phase transition
We discuss the compatibility of the degenerate scalar scenario and EWPT qualitatively. As mentioned in Sec. III, the orthogonality of the mixing matrix (32) is important for the degenerate scalar scenario. Here, using the elements of the mixing matrix, the mixing term for SU(2)L doublet-singlet can be expressed as
$\delta_{1}$ $=\frac{2}{v_{1}v_{s}}\left(O_{1i}O_{3i}m_{H_{i}}^{2}\right),$ (65)
$\delta_{2}$ $=\frac{2}{v_{2}v_{s}}\left(O_{2i}O_{3i}m_{H_{i}}^{2}\right).$ (66)
Compared to Eqs. (50) and (53), it can be seen that the DM-quark scattering is controlled by $\delta_{1}$ and $\delta_{2}$. In other words, the degenerate scalar scenario is realized by $\delta_{1}$ and $\delta_{2}$.
Strong first-order EWPT is essential for EWBG, and the condition is Arnold and McLerran (1987); Bochkarev and Shaposhnikov (1987); Funakubo and Senaha (2009);
$\frac{v_{C}}{T_{C}}\gtrsim 1,$ (67)
where $T_{C}$ is the critical temperature at which the effective potential has two degenerate minima, and $v_{C}$ is the nonzero Higgs VEV at $T=T_{C}$. Possible origins of first-order EWPT are the finite-temperature boson loop and the structure of the tree-level potential. In the former case, the addition of multiple bosons is effective, and the 2HDM specializes in this type of EWPT. On the other hand, in the latter case, the existence of a single scalar is important, and the CxSM causes this type of EWPT. Since the 2HDMS has multiple bosons and a singlet scalar that does not exist in the SM, EWPT of both types can occur. First, we qualitatively analyze tree-level driven EWPT using the so-called high temperature (HT) potential, which is composed of the tree-level potential and thermal masses.
EWPT pattern of 2HDMS is $\quantity(\langle\Phi_{1}\rangle,\langle\Phi_{2}\rangle,\langle S\rangle)=(0,0,v_{S}^{\prime})\to(v_{1},v_{2},v_{S})$, where $\langle\Phi_{i}\rangle$ and $\langle S\rangle$ are defined as the VEVs of the doublet fields $\Phi_{i}$ and the singlet field $S$, respectively. To discuss EWPT, we take the Landau gauge and denote the classical background fields of the Higgs doublets and singlet as
$\langle\Phi_{i}\rangle=\left(\begin{array}[]{ll}0&\varphi_{i}\end{array}\right)^{\top}/\sqrt{2},\quad\langle S\rangle=\varphi_{S}/\sqrt{2}.$ (69)
The tree-level potential (1) can be rewritten as
$V_{0}\left(\varphi_{1},\varphi_{2},\varphi_{S}\right)$ $=\frac{m_{1}^{2}}{2}\varphi_{1}^{2}+\frac{m_{2}^{2}}{2}\varphi_{2}^{2}-m_{3}^{2}\varphi_{1}\varphi_{2}+\frac{\lambda_{1}}{8}\varphi_{1}^{4}+\frac{\lambda_{2}}{8}\varphi_{2}^{4}+\frac{\lambda_{345}}{4}\left(\varphi_{1}\varphi_{2}\right)^{2}$
$+\frac{\delta_{1}}{8}\varphi_{1}^{2}\varphi_{S}^{2}+\frac{\delta_{2}}{8}\varphi_{2}^{2}\varphi_{S}^{2}+\frac{b_{2}}{4}\varphi_{S}^{2}+\frac{d_{2}}{16}\varphi_{S}^{4}+\sqrt{2}a_{1}\varphi_{S}+\frac{b_{1}}{4}\varphi_{S}^{2}.$ (70)
The HT potential is defined as
$V^{\mathrm{HT}}\left(\varphi_{1},\varphi_{2},\varphi_{S};T\right)=V_{0}\left(\varphi_{1},\varphi_{2},\varphi_{S}\right)+\left(\Sigma_{h1}\varphi_{1}^{2}+\Sigma_{h2}\varphi_{2}^{2}+\Sigma_{S}\varphi_{S}^{2}\right)T^{2},$ (71)
where
$\Sigma_{h1}$ $=\frac{\delta_{1}}{48}+\frac{g_{1}^{2}}{32}+\frac{3g_{2}^{2}}{32}+\frac{\lambda_{1}}{8}+\frac{\lambda_{3}}{12}+\frac{\lambda_{4}}{24}+\frac{y_{t}^{2}}{8}+\frac{y_{b}^{2}}{8},$
$\Sigma_{h2}$ $=\frac{\delta_{2}}{48}+\frac{g_{1}^{2}}{32}+\frac{3g_{2}^{2}}{32}+\frac{\lambda_{2}}{8}+\frac{\lambda_{3}}{12}+\frac{\lambda_{4}}{24}+\frac{y_{t}^{2}}{8}+\frac{y_{b}^{2}}{8},$
$\Sigma_{S}$ $=\frac{d_{2}}{24}+\frac{\delta_{1}}{24}+\frac{\delta_{2}}{24},$ (72)
with $g_{2}$, $g_{1}$, $y_{t}$ and $y_{b}$ being SU(2)L , U(1)Y, top Yukawa and bottom Yukawa couplings. It should be emphasized that the thermal cubic term of the field is necessary for loop-driven EWPT. Since the HT potential does not include that term, it is optimal for analyzing EWPT derived from the tree-level potential.
For convenience, $\varphi_{i}$ and $\varphi_{S}$ are parameterized as
$\varphi_{1}=z\cos\theta\cos\beta,\quad\varphi_{2}=z\cos\theta\sin\beta,\quad\varphi_{S}=z\sin\theta+\tilde{v}_{S},$ (73)
where $\tilde{v}_{S}$ denotes the minimum on the $\varphi_{S}$ axis, which is always nonzero because of $a_{1}=0$ in our model. The HT potential (71) becomes
$V(z,\theta,\beta;T)=c_{0}+c_{1}z+\left(c_{2}+c_{2}^{\prime}T^{2}\right)z^{2}-c_{3}z^{3}+c_{4}z^{4},$ (74)
and at $T=T_{C}$, the above potential takes the form
$V\left(z,\theta,\beta;T_{C}\right)=c_{4}z^{2}\left(z-z_{C}\right)^{2},\quad z_{C}=\frac{c_{3}}{2c_{4}},$ (75)
where
$c_{3}$ $=\frac{1}{4}\tilde{v}_{S}\sin\theta\left(\delta_{1}\cos^{2}\beta\cos^{2}\theta+\delta_{2}\cos^{2}\theta\sin^{2}\beta+d_{2}\sin^{2}\theta\right),$ (76)
$c_{4}$ $=\frac{1}{16}\left(2\lambda_{1}\cos^{4}\beta\cos^{4}\theta+2\lambda_{2}\cos^{4}\theta\sin^{4}\beta+2\delta_{2}\cos^{2}\theta\sin^{2}\beta\sin^{2}\theta\right.$
$\left.+d_{2}\sin^{4}\theta+2\cos^{2}\beta\left(2\lambda_{345}\cos^{4}\theta\sin^{2}\beta+\delta_{1}\cos^{2}\theta\sin^{2}\theta\right)\right).$ (77)
From this potential, $v_{C}$ and $T_{C}$ are derived, which is
$v_{C}$ $=\sqrt{\frac{{v}_{SC}^{\prime}({v}_{SC}^{\prime}-{v}_{SC})\left(\delta_{1}\cos{\beta_{C}}^{2}+\delta_{2}\sin{\beta_{C}}^{2}\right)}{\lambda_{1}\cos_{\beta_{C}}^{4}+\lambda_{2}\sin_{\beta_{C}}^{4}+2\lambda_{345}\sin_{\beta_{C}}^{2}\cos_{\beta_{C}}^{2}}}$ (78)
$T_{C}$ $=\sqrt{\frac{-m_{1}^{2}\cos^{2}{\beta}-m_{2}^{2}\sin^{2}{\beta}+2m_{3}^{2}\cos{\beta}\sin{\beta}-\frac{1}{4}{v}_{SC}^{\prime}(\delta_{1}\cos^{2}{\beta}+\delta_{2}\sin^{2}{\beta})}{\Sigma_{h1}\cos^{2}{\beta}+\Sigma_{h2}\sin^{2}{\beta}}}.$ (79)
The physical quantities evaluated by $T_{C}$ are labeled with the subscript $C$. We can find that the condition of strong first-order EWPT (67) requires large $\delta_{1}$ and $\delta_{2}$.
However, as mentioned above, small $\delta_{1},\delta_{2}$ are essential for realizing the degenerate scalar scenario. Therefore, we do not entrust EWPT to the tree-level double-single mixing $\delta_{1},\delta_{2}$, but to the thermal loop. The 1-loop effective potential can be written as follows:
$V_{\text{eff}}\left(\varphi_{1},\varphi_{2},\varphi_{S};T\right)$ $=V_{0}(\varphi_{1},\varphi_{2},\varphi_{S})+V_{1}(\varphi_{1},\varphi_{2},\varphi_{S};T),$
$=V_{0}(\varphi_{1},\varphi_{2},\varphi_{S})+\sum_{i}n_{i}\left[V_{\mathrm{CW}}\left(\bar{m}_{i}^{2}\right)+\frac{T^{4}}{2\pi^{2}}I_{B,F}\left(\frac{\bar{m}_{i}^{2}}{T^{2}}\right)\right],$ (80)
where $T$ represents temperature. The subscript $i$ represents the particles contained in the model, and $n_{i}$ is the degree of freedom of each particle. The effective potentials at zero temperature and at the finite temperature are given by Weinberg (1973); Jackiw (1974); Dolan and Jackiw (1974)
$V_{\mathrm{CW}}\left(\bar{m}_{i}^{2}\right)$ $=\frac{\bar{m}_{i}^{4}}{64\pi^{2}}\left(\ln\frac{\bar{m}_{i}^{2}}{\bar{\mu}^{2}}-c_{i}\right),$ (81)
[…]
where $c_{i}=3/2$ for scalars and fermions and $c_{i}=5/6$ for gauge bosons. $\bar{m_{i}}$ is the field-dependent masses of each particle $i$. $I_{B}$ with the minus sign represents the boson contribution, while $I_{F}$ with the plus sign represents the fermion one. $\bar{\mu}$ is a renormalization scale.
In the following calculations, the condition (67) is checked using the 1-loop effective potential (80). In order to prevent the breakdown of perturbation theory due to boson multi-loop, the Parwani resummation, in which all the Matsubara frequency modes are resummed, is used Parwani (1992). Specifically, field-dependent masses $\bar{m}_{i}^{2}$ in Eq. (80) are replaced by thermally corrected ones.
V Numerical results
Before giving benchmark points, we mention the theoretical and experimental constraints imposed on the parameters. A requirement on the scalar potential that is bounded from below is given by Nie and Sher (1999); Kanemura et al. (1999)
$\lambda_{1}>0,~{}\lambda_{2}>0,~{}\lambda_{3}+\lambda_{4}-\lambda_{5}>-\sqrt{\lambda_{1}\lambda_{2}},~{}d_{2}>0.$ (83)
The quartic couplings should also satisfy the following conditions from the tree-level unitarity Chen et al. (2015); Akeroyd et al. (2000)
$\lambda_{1}<\frac{8\pi}{3},~{}\lambda_{2}<\frac{8\pi}{3},~{}\lambda_{345}<{8\pi},~{}\delta_{1}<16\pi,~{}\delta_{2}<16\pi,~{}d_{2}<\frac{16\pi}{3}.$ (84)
The constraint from the perturbativity is as follows Aoki et al. (2022):
$\left|\lambda_{j}\right|,d_{2}<4\pi~{}(j=1-5).$ (85)
Furthermore, the sizes of $\delta_{1}$ and $\delta_{2}$ could be constrained by a global minimum condition at zero temperature Chen et al. (2020). The energy of the electroweak vacuum has to be lower than that of the local vacuum on the $\varphi_{S}$ axis. Experimental constraints here are primarily related to electroweak precision data and flavor experiments. By assuming a mass degeneracy between the charged scalar boson $H^{\pm}$ and the neutral scalar boson $H$ or the CP-odd scalar boson $A$, i.e., $m_{H^{\pm}}\simeq m_{H}\text{ or }m_{A}$, the electroweak precision data is satisfied Haber and O’Neil (2011). In addition, for Type-I 2HDMS, $B_{d}\rightarrow\mu\mu$ process requires $\tan{\beta}\gtrsim 1.75$ for $m_{H^{\pm}}=500$ GeV Haller et al. (2018). We should emphasize that the Higgs coupling measurement usually places a constraint on $\cos(\beta-\alpha)$, but not now since our analysis is based on the degenerate scalar scenario.
Inputs $v$ $v_{S}$ $m_{H_{1}}$ $m_{H_{2}}$ $m_{H_{3}}$ $m_{H^{\pm}}$ $m_{A}$ $m_{\chi}$ $\alpha_{1}$ $\alpha_{2}$ $\alpha_{3}$ $\tan{\beta}$ ) $a_{1}$  $m_{3}^{2}$
BP1 246.22 246.22 125.0 124.5 124.0 500 500 200 $\pi/4$ $\pi/4$ 0.01 2.0 -6550 10
BP2 246.22 200 125.0 124.5 124.0 500 500 200 $\pi/4$ $\pi/4$ 0.01 2.0 -700000 10
BP3 246.22 30 125.0 124.5 124.0 500 500 200 $\pi/4$ $\pi/4$ 0.01 2.0 -323000 10
Outputs $\delta_{1}$ $\delta_{2}$ $d_{2}$ $\lambda_{1}$ $\lambda_{2}$ $\lambda_{3}$ $\lambda_{4}$ $\lambda_{5}$ $m_{1}^{2}$ $m_{2}^{2}$ $b_{1}$ $b_{2}$ $v_{1}$ $v_{2}$
BP1 0.0069 0.0017 0.51 1.27 0.32 8.25 -4.12 -4.12 -7863 -7812 -39962 24523 110 220
BP2 0.0085 0.0020 0.53 1.27 0.32 8.25 -4.12 -4.12 -7843 -7807 -35050 34328 110 220
BP3 0.057 0.014 0.54 1.27 0.32 8.25 -4.12 -4.12 -7771 -7790 -24774 54308 110 220
Table 2: Input and output parameters.
The input and output parameters that satisfy the above constraints are summarized in Table 2. The masses of three Higgs bosons are degenerate within 1 GeV. In addition, these three benchmark points (BPs) reflect the EWPT discussion in Sec. IV. The difference between input parameters among BPs is the size of $v_{S}$. The parameter $a_{1}$ is changed to keep the value of $d_{2}$ perturbatively appropriate.
$v_{1C}$ [GeV] $v_{2C}$ [GeV] $v_{SC}$ [GeV] $v_{SC}^{\prime}$ [GeV] $v_{C}/T_{C}$
 BP1 105.8 213.2 245.7 246.3 $\frac{238.0}{69.6}=3.4$
 BP2 105.8 213.2 199.5 200.1 $\frac{238.0}{69.6}=3.4$
 BP3 105.7 213.2 30.0 30.6 $\frac{238.0}{69.6}=3.4$
Table 3: $T_{C}$ and corresponding VEVs for three BPs in Table. 2. Here, the calculations are performed by cosmoTransitions Wainwright (2012).
The public code cosmoTransitions Wainwright (2012) is used to evaluate EWPT. In Table 3, $T_{C}$ and corresponding VEVs for BPs are shown, and strong first-order EWPT occurs in all BPs. The effect of $\lambda_{i}$ is known to be significant in the thermal loop-derived EWPT, and since the value of $\lambda_{i}$ is the same for all BPs, the Higgs VEVs and $T_{C}$ are almost identical in three cases. In addition, the small change in the singlet VEV before and after EWPT shows that the tree-level effect by the singlet introduction does not cause EWPT this time. In fact, when using the HT potential (71), EWPT becomes second-order.
Figure 2: Spin-independent cross section for the scattering between the DM and nucleus $\sigma_{\mathrm{SI}}$ as a function of $m_{H_{2}}$.
Next, we show the results of numerical calculations related to the DM using the public code micrOMEGAs Belanger et al. (2010, 2021). In Fig. 2, the spin-independent cross section for the scattering between the DM and nucleus $\sigma_{\mathrm{SI}}$ as a function of $m_{H_{2}}$ is shown. Here, $m_{H_{2}}$ is used once as a variable, and the other parameters are BPs in Table 2. The shaded regions are excluded by the LZ experiment. For all BPs, there is a large dip around $m_{H_{2}}\simeq 125$ GeV. This indicates that the degenerate scalar scenario is important for the suppression of DM and nucleon scattering. Furthermore, BP1 has the smallest $\delta_{1}$ and $\delta_{2}$, which describes the SU(2)L doublet-singlet mixing, and thus satisfies the experimental results outside of the degenerate region. Although the elastic neutrino-nucleon scattering makes DM detection difficult in some regions, this does not rule out the existence of the WIMP-DM O’Hare (2021).
Figure 3: (Left panel) DM relic density $\Omega_{\chi}h^{2}$ as a function of DM mass $m_{c}hi$ for the BPs in Table. 2 is shown. The black dashed line represents the observed value (86), and the region above this corresponds to overproduction. (Right panel) scaled scattering cross section between the DM $\chi$ and quarks $\tilde{\sigma}_{\mathrm{SI}}$ as a function of DM mass $m_{\chi}$ for all BPs is shown. The dotted line represents the bound of the LZ experiment, and the region above this is excluded. The red dots on both panels indicate where the observed relic density and the predicted one from the 2HDMS are equal.
The left panel of Fig. 3 shows $\Omega_{\chi}h^{2}$ as a function of $m_{\chi}$ for all BPs in Table. 2. The value of the DM relic density $\Omega_{\mathrm{DM}}h^{2}$ must not exceed the observed value Aghanim et al. (2020):
$\Omega_{\mathrm{DM}}h^{2}=0.1200\pm 0.0012,$ (86)
which corresponds to the black dashed line in the graph. The red dots indicate the points where the calculated value equals the observed value (86).
On the other hand, the right panel of Fig. 3 shows the $m_{\chi}$ dependence of the spin-independent scattering cross section between DM and nucleus. For $\Omega_{\chi}<\Omega_{\mathrm{DM}}$, we should scale $\sigma_{\mathrm{SI}}$ as
$\tilde{\sigma}_{\mathrm{SI}}=\left(\frac{\Omega_{\chi}}{\Omega_{\mathrm{DM}}}\right)\sigma_{\mathrm{SI}}.$ (87)
Here, the scaled cross section is shown for each BP. The dotted line is the bound of the LZ experiment, and the region above this is excluded. The red dots correspond to those of the left panel. It can be seen that many DM mass regions are still allowed, and the DM $\chi$ may carry all of the existing DM and explain the current experimental results.
VI Summary and discussion
In this paper, we have investigated the compatibility of the degenerate scalar scenario and strong first-order EWPT in the 2HDMS, which contains two Higgs doublets and one singlet scalar. In this model, the imaginary part of $S$ behaves as a DM $\chi$, and CP-even scalars mix to form three Higgs particles $H_{i}~{}(i=1,2,3)$. The scattering of DM and quarks is mediated by $H_{i}$, and we have found that this scattering is suppressed in the region where Higgs masses are degenerate, as shown in previous studies on pNG DM Abe et al. (2021). The suppression of SU(2)L doublet-singlet mixing terms $\delta_{1}$ and $\delta_{2}$, which are related to the orthogonality of the mixing matrix that transforms the current eigenstates to the mass eigenstates of the CP-even scalars, is important for the degenerate scalar scenario.
Strong first-order EWPT, an essential element of EWBG, can be derived from two sources in the effective potential: the thermal loop and the tree-level structure. Since the tree-level driven EWPT is induced by the presence of the singlet field, i.e., large $\delta_{1},~{}\delta_{2}$ are important, it was pointed out here that this is inconsistent with the degenerate scalar scenario (the discussion in the CxSM is given in ref. Cho et al. (2021)). Therefore, as is common with the 2HDM, it is best to leave strong first-order EWPT to the loop effect.
We have set the benchmark points based on the qualitative discussion of EWPT and confirmed that strong first-order EWPT actually occurs. Since we focus on the degenerate scalar scenario, the DM-quark spin-independent cross section is suppressed even in regions that satisfy the DM observed relic density, and we were able to explain the constraint from the LZ experiment in many regions. Therefore, it was found that current observational data and experimental results can be explained by the DM $\chi$ in this model and that it is also consistent with the condition given by strong first-order EWPT.
$\lambda_{111}/\lambda_{hhh}^{\mathrm{SM}}$ $\lambda_{112}/\lambda_{hhh}^{\mathrm{SM}}$ $\lambda_{113}/\lambda_{hhh}^{\mathrm{SM}}$ $\Delta\lambda_{11}^{\mathrm{total}}$
BP1 $0.799$ $-0.00872$ $0.571$ 0.362
BP2 $0.780$ $0.0109$ $0.544$ 0.335
BP3 $0.701$ $0.0921$ $0.431$ 0.224
Table 4: The deviation from the SM values for the Higgs trilinear couplings $\lambda_{111}$, $\lambda_{112}$, $\lambda_{113}$ and their sum defined as Eq. (89).
Finally, we briefly discuss the possibility of verifying the model in future collider experiments. A distinctive feature of the thermal loop-derived EWPT is a shift of the Higgs trilinear coupling
$\mathcal{L}$ $=\lambda_{ijk}H_{i}H_{j}H_{k}~{}~{}~{}(i,j,k=1,2,3),$ (88)
from the SM prediction Kanemura et al. (2005); Grojean et al. (2005). For example, we focus on the pair production of $H_{1}$ through off-shell CP-even scalars $H_{i}$, i.e., $H_{i}\to H_{1}H_{1}$. We compare each trilinear coupling $\lambda_{11i}~{}(i=1,2,3)$ in three benchmark points (BP1, BP2, BP3) with the SM. Furthermore, since amplitudes mediated by off-shell scalars $H_{i}$ interfere with each other, we also compare the sum of trilinear couplings $\lambda_{111},\lambda_{112}$ and $\lambda_{113}$ to the SM prediction. The results are shown in Table. 4, where $\Delta\lambda_{11}^{\mathrm{total}}$ is defined as
$\Delta\lambda_{11}^{\mathrm{total}}=\frac{(\lambda_{111}+\lambda_{112}+\lambda_{113})-\lambda^{\mathrm{SM}}_{hhh}}{\lambda^{\mathrm{SM}}_{hhh}},$ (89)
The Higgs trilinear coupling of this model is expected to differ from that of SM by about 30%, and it may be confirmed at the High-Energy Large Hadron Collider (HE-LHC) and/or the International Linear Collider (ILC), whose accuracy is estimated by ref. Shiltsev (2019). We leave the detailed analysis and the model verification in other experiments, such as the search for the degenerate Higgs at the lepton collider and observation of gravitational waves from EWPT, to future research.
Acknowledgements.
We are grateful to Ayana Okuma and Chiaki Nose for their valuable discussions. The work of GCC is supported by JSPS KAKENHI Grant No. 22K03616.
Appendix A Input and output parameters
The following shows the relationship between the input and output parameters. However, $m_{3}^{2}$ and $a_{1}$ remain as input parameters. The index $i$ runs 1 to 3, and the sum symbol $\quantity(\sum_{i})$ is omitted.
$m_{1}^{2}$ $=\frac{m_{3}^{2}v_{2}}{v_{1}}-\frac{\lambda_{1}v_{1}^{2}}{2}-\frac{\lambda_{345}v_{2}^{2}}{2}-\frac{\delta_{1}v_{s}^{2}}{4},$ (90)
$m_{2}^{2}$ $=\frac{m_{3}^{2}v_{1}}{v_{2}}-\frac{\lambda_{2}v_{2}^{2}}{2}-\frac{\lambda_{345}v_{1}^{2}}{2}-\frac{\delta_{2}v_{s}^{2}}{4},$ (91)
$\lambda_{1}$ $=\frac{1}{v_{1}^{2}}\left(O_{1i}^{2}m_{H_{i}}^{2}-\frac{m_{3}^{2}v_{2}}{v_{1}}\right),$ (92)
$\lambda_{2}$ $=\frac{1}{v_{2}^{2}}\left(O_{2i}^{2}m_{H_{i}}^{2}-\frac{m_{3}^{2}v_{1}}{v_{2}}\right),$ (93)
$\lambda_{3}$ $=\frac{1}{v_{1}v_{2}}\left(O_{1i}O_{2i}m_{H_{i}}^{2}+m_{3}^{2}\right)-\lambda_{4}-\lambda_{5},$ (94)
$\lambda_{4}$ $=\frac{2}{v^{2}}\left(\frac{m_{3}^{2}}{\sin\beta\cos\beta}-m_{H^{\pm}}^{2}\right)-\lambda_{5},$ (95)
$\lambda_{5}$ $=\frac{1}{v^{2}}\left(\frac{m_{3}^{2}}{\sin\beta\cos\beta}-m_{A}^{2}\right),$ (96)
$\delta_{1}$ $=\frac{2}{v_{1}v_{s}}\left(O_{1i}O_{3i}m_{H_{i}}^{2}\right),$ (97)
$\delta_{2}$ $=\frac{2}{v_{2}v_{s}}\left(O_{2i}O_{3i}m_{H_{i}}^{2}\right),$ (98)
$d_{2}$ $=\frac{2}{v_{s}^{2}}\left(\frac{\sqrt{2}a_{1}}{v_{s}}+O_{3i}^{2}m_{H_{i}}^{2}\right),$ (99)
$b_{2}$ $=\frac{-4\sqrt{2}a_{1}-2b_{1}v_{s}-\delta_{1}v_{1}^{2}v_{s}-\delta_{2}v_{2}^{2}v_{s}-d_{2}v_{s}^{3}}{2v_{s}},$ (100)
$b_{1}$ $=-m_{\chi}^{2}-\frac{\sqrt{2}a_{1}}{v_{s}}.$ (101)